OpenStudy (anonymous):

two automobiles set out simultaneously from places 240 kilometers apart. they traveled toward each other and passed at the end of 2 hours. the speed of one automobile is 10 kph greater than the other. find the speed of each.

OpenStudy (phi):

a bit tricky, as you have 2 equations and 2 unknowns. we know speed * time= distance call the speed of car 1= s and the speed of car 2= s+10 (it's 10 kph faster than car 1) call the distance car 1 travels= d call the distance car 2 travels = 240 -d (between the 2 cars, we go 240 km) for car 1, after traveling for 2 hours: 2*s= d also, we know time= distance/speed and time = 2 hours for both cars d/s = (240-d)/(s+10) also d=2s so substitute 2s for d in the 1st equation 2s/s = (240-2s)/(s+10) solve for s